Find the projection on the x-axis if AB = 10 and theta = 30 degrees.
Answer by venugopalramana(3286)

PROJECTION ON X AXIS OF AB = AB*COS(THETA)...WHERE THETA IS ANGLE BETWEEN X AXIA AND LINE AB.
SO PROJECTION =10*COS(30)=10/2=5